# Shuchang Zhang, Hui Zhang, Hongxia Wang

### v5

@@ 4.1 A toy example
The corresponding saddle-point problem is $\min_{x\in\mathbb{R}}\max_{y\in\mathbb{R}}\{xy+f(x)-g^{*}(y)\}$, where $g^{*}(y)=\max_{x}[yx-\max\{1-x,0\}]=y+\delta_{[-1,0]}(y)$ and $\delta_{[-1,0]}$ is the indicator function of the interval $[-1,0]$. We denote the optimal set $X^{*}=[1,+\infty)=\arg\min_{x\mathbb{\in\mathbb{R}}}f(x)+g(x)$ and the primal-dual objective function $F(x,y)=xy+f(x)-g^{*}(y)$. Let us solve the saddle-point set $\{(x^{*},y^{*}):F(x^{*},y)\leq F(x^{*},y^{*})\leq F(x,y^{*})\text{ for all }(x,y)\in\mathbb{R}^{2}\}$ of $F(x,y)$. Fix $x^{*}\geq 1$; then
@@
If $x^{*}=1$, assume that $-1\leq y^{*}<0$, then $F(1,y^{*})=0$, while $F(x,y^{*})=y^{*}(x-1)+\max\{-x,0\}$, and there exists $x=2$ such that $F(2,y^{*})=y^{*}<F(1,y^{*})$, which leads to a contradiction. Therefore, the saddle-point set is $\Omega=\{(x^{*},y^{*}):x^{*}\geq 1,y^{*}=0\}$.
@@ 4.2 A 2D toy example
We consider a $\frac{1}{\gamma}$-contractive operator $T_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ from [41] for any $\mathbf{x}=(x_{1},x_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$,
@@ 4.3 An infinite-dimensional example
We consider an infinite-dimensional problem on $\ell^{2}$ from [7], the space of square summable sequences. For any $\mathbf{x}=(x_{0},x_{1},x_{2},\ldots)\in\ell^{2}$, the norm is $\left\|\mathbf{x}\right\|=\left(\sum_{k\in\mathbb{N}}\left|x_{i}\right|^{2}\right)^{\frac{1}{2}}$. We consider the following operator inclusion problem:
@@
where $\mathcal{S}(\mathbf{x})=(0,x_{0},x_{1},\ldots)$ is the right-shift operator and $\mathbf{b}=(1,-1,0,\ldots)$. Since $\mathcal{A}_{1},\mathcal{A}_{2}$ are nonexpansive, $\mathcal{A}_{1},\mathcal{A}_{2}$ are maximal monotone [2, Example 20.29]. Additionally, by solving the above inclusion problem (22), we obtain $\mathbf{x}^{*}=(1,0,\ldots)$. Using $\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&I\end{pmatrix}$ and positive semidefinite preconditioner $\mathcal{M}=\begin{pmatrix}2I&-I\\ -I&\frac{1}{2}I\end{pmatrix}$, we compare the PPP iteration with $\lambda_{k}=0.5,1,1.8$. We set $N=10^{3},\tau=0.5,s=2$ and initialize $\mathbf{x}^{0}=(0.9,1,0,\ldots)$ for HPPP and PPP, and the anchor point is the same as the initial point, i.e., $\mathbf{a}=(\mathbf{x}^{0},\mathbf{x}^{0})$. For this example, we present visual results about the residual and the distance to the solution. As illustrated in Figure 5, HPPP outperforms PPP in terms of both $\left\|\mathbf{x}^{k}-\mathbf{x}^{*}\right\|$ and $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$.
@@ 4.4 Image deblurring
First, we compared CP [16], PPP [9], and HPPP for TV-regularized IR problems. For TV-regularized IR problems,
@@
Second, we compared RED and RED-PRO. Following [47, 20], a $9\times 9$ uniform point spread function (PSF) and a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images. We finally obtained the degraded images with the noise level $\sigma=\sqrt{2}$. The original RGB image is converted into the YCbCr image, all algorithms are applied to the luminance channel, and then the reconstruction image is returned to RGB space to obtain the final image. PSNR is measured on the luminance channel of the ground truth and the restored images. Table 3 shows the PSNR values of the restoration results on RED, RED-PRO, RRP, GraRED-P3, and GraRED-HP3. From the deblurring experiment of Table 3 and Figure 9, GraRED-P3 and GraRED-HP3 achieve better performance than RED, RED-PRO, and RRP, which illustrates that KM or Halpern iteration used in PPP methods is effective. We further compare the differences among GraRED-P3, GraRED-HP3, and restarted HPPP under six different settings with $\lambda=40,\tau=s=1,\lambda_{k}=0.2,\mu_{k}=\frac{1}{k+1},N=500,q=100$. In Table 5, we report the average PSNR values of 15 test images using six different blur kernels. When the anchor point and the initial point coincide, GraRED-HP3 achieves the best performance on some IR tasks. Moreover, by dynamically updating the anchor points, restarted HPPP further boosts the performance of GraRED-HP3.
@@
Additionally, we trained a FNE-DnCNN ($D_{1}$) using real spectral normalization [49], which is employed in RED and RED-PRO. We also compared our results with state-of-the-art methods using advanced denoisers, such as DRUNet ($D_{2}$[65] and GS ($D_{3}$[32, 33]. The PnP-DYS method [59] uses the Davis-Yin splitting (DYS) algorithm. For GraRED-HP3, the anchor point is the same as the initial point. The default parameters are presented in Table 4. As shown in Table 6, Restarted HPPP with advanced denoisers achieves competitive or even superior performance compared to state-of-the-art methods. In Figure 8, we compare GraRED-P3 ($\lambda_{k}=0.1$) and GraRED-HP3 ($\mathbf{a}=\mathbf{u}^{0},\mu_{k}=\frac{1}{k+1}$). The results demonstrate that GraRED-HP3 achieves better recovery performance and exhibits a faster fixed-point convergence rate than GraRED-P3.
Table 3: Recovery results are obtained by different algorithms with trainable nonlinear reaction-diffusion (TNRD) denoiser [18]. The best two results are highlighted in red and blue, respectively.
@@
RED [47] 27.19 26.00 26.46 29.21 30.16 32.41 28.46 28.36 29.58
RED-PRO [20] 27.43 26.05 26.39 28.99 29.98 32.41 28.25 28.21 29.38
DPIR [65] 28.47 27.02 27.63 30.24 30.89 33.56 29.62 30.43 30.54
PnP-DRS [33] 29.79 26.92 27.46 29.70 29.59 33.25 29.94 30.66 29.46
PnP-PGD [32] 27.18 26.28 26.73 29.17 29.14 32.84 29.14 29.87 30.37
PnP-DYS [59] 27.22 26.27 26.62 28.97 29.04 32.87 29.06 29.88 28.68
@@ 4.5 Image inpainting
In this section, we use the proposed algorithm to solve TV-regularized image inpainting problems and compare their numerical results with CP [16], PPP [9], and HPPP algorithms. The discrete image inpainting model is
@@
Both GraRED-P3 and GraRED-HP3 use DnCNN [49], other parameters are used:
@@ 5 Conclusions
This paper introduces a Halpern-type PPP (HPPP) algorithm, which has the advantage of strong convergence over PPP. In addition, GraRED-HP3 is proposed by integrating HPPP with denoiser priors for IR problems, which is an accelerated PnP-ADMM. Advantages of HPPP are demonstrated through several toy examples, and GraRED-HP3 achieves state-of-the-art performance on IR experiments. In the future, we plan to extend the definition of $\mathcal{M}$-monotonicity to $\mathcal{M}$-comonotonicity for nonconvex case [4].
@@ Acknowledgments
The authors thank Prof. Defeng Sun for bringing the interesting work [52] to our attention and Mr. Feng Xue for his insightful discussions. We sincerely appreciate Associate Editor Michael Elad and the anonymous referees for their thorough review and valuable comments, which have substantially improved this paper.
@@ Appendix A Boundedness
First, we study the boundedness and asymptotic regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by (5). To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce an important lemma from [30, Lemma 2.5].
@@
Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ [9, Proposition 2.3]. Denote
@@
First, we prove $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ is bounded. For any $\mathbf{u}^{*}\in\mathrm{Fix}(T)$, we have
@@ Appendix B $\mathcal{M}$-projection
As mentioned in [3, 15], we can introduce the following notion of $\mathcal{M}$-projection.

### v6

@@ 4.1 A toy example
The corresponding saddle-point problem is $\min_{x\in\mathbb{R}}\max_{y\in\mathbb{R}}\{xy+f(x)-g^{*}(y)\}$, where $g^{*}(y)=\max_{x}[yx-\max\{1-x,0\}]=y+\delta_{[-1,0]}(y)$ and $\delta_{[-1,0]}$ is the indicator function of the interval [1,0][-1,0]. We denote the optimal set $X^{*}=[1,+\infty)=\arg\min_{x\mathbb{\in\mathbb{R}}}f(x)+g(x)$ and the primal-dual objective function $F(x,y)=xy+f(x)-g^{*}(y)$. Let us solve the saddle-point set $\{(x^{*},y^{*}):F(x^{*},y)\leq F(x^{*},y^{*})\leq F(x,y^{*})\text{ for all }(x,y)\in\mathbb{R}^{2}\}$ of $F(x,y)$. Fix $x^{*}\geq 1$; then
@@
If $x^{*}=1$, assume that $1\leq y^{*}<0$, then $F(1,y^{*})=0$, while $F(x,y^{*})=y^{*}(x-1)+\max\{-x,0\}$, and there exists $x=2$ such that $F(2,y^{*})=y^{*}<F(1,y^{*})$, which leads to a contradiction. Therefore, the saddle-point set is $\Omega=\{(x^{*},y^{*}):x^{*}\geq 1,y^{*}=0\}$.
@@ 4.2 A 2D toy example
We consider a $\frac{1}{\gamma}$-contractive operator $T_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ from [Park2022] for any $\mathbf{x}=(x_{1},x_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$,
@@ 4.3 An infinite-dimensional example
We consider an infinite-dimensional problem on $\ell^{2}$ from [bc24], the space of square summable sequences. For any $\mathbf{x}=(x_{0},x_{1},x_{2},\ldots)\in\ell^{2}$, the norm is $\left\|\mathbf{x}\right\|=\left(\sum_{k\in\mathbb{N}}\left|x_{i}\right|^{2}\right)^{\frac{1}{2}}$. We consider the following operator inclusion problem:
@@
where $\mathcal{S}(\mathbf{x})=(0,x_{0},x_{1},\ldots)$ is the right-shift operator and $\mathbf{b}=(1,-1,0,\ldots)$. Since $\mathcal{A}_{1},\mathcal{A}_{2}$ are nonexpansive, $\mathcal{A}_{1},\mathcal{A}_{2}$ are maximal monotone [BCombettes, Example 20.29]. Additionally, by solving the above inclusion problem (22), we obtain $\mathbf{x}^{*}=(1,0,\ldots)$. Using $\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&I\end{pmatrix}$ and positive semidefinite preconditioner $\mathcal{M}=\begin{pmatrix}2I&-I\\ -I&\frac{1}{2}I\end{pmatrix}$, we compare the PPP iteration with $\lambda_{k}=0.5,1,1.8$. We set $N=10^{3},\tau=0.5,s=2$ and initialize $\mathbf{x}^{0}=(0.9,1,0,\ldots)$ for HPPP and PPP, and the anchor point is the same as the initial point, i.e., $\mathbf{a}=(\mathbf{x}^{0},\mathbf{x}^{0})$. For this example, we present visual results about the residual and the distance to the solution. As illustrated in Figure 5, HPPP outperforms PPP in terms of both $\left\|\mathbf{x}^{k}-\mathbf{x}^{*}\right\|$ and $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$.
@@ 4.4 Image deblurring
First, we compared CP [Chambolle2011], PPP [Bredies2022], and HPPP for TV-regularized IR problems. For TV-regularized IR problems,
@@
Second, we compared RED and RED-PRO. Following [Romano2017, cohen2021regularization], a $9\times 9$ uniform point spread function (PSF) and a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images. We finally obtained the degraded images with the noise level $\sigma=\sqrt{2}$. The original RGB image is converted into the YCbCr image, all algorithms are applied to the luminance channel, and then the reconstruction image is returned to RGB space to obtain the final image. PSNR is measured on the luminance channel of the ground truth and the restored images. Table 3 shows the PSNR values of the restoration results on RED, RED-PRO, RRP, GraRED-P3, and GraRED-HP3. From the deblurring experiment of Table 3 and Figure 9, GraRED-P3 and GraRED-HP3 achieve better performance than RED, RED-PRO, and RRP, which illustrates that KM or Halpern iteration used in PPP methods is effective. We further compare the differences among GraRED-P3, GraRED-HP3, and restarted HPPP under six different settings with $\lambda=40,\tau=s=1,\lambda_{k}=0.2,\mu_{k}=\frac{1}{k+1},N=500,q=100$. In Table 5, we report the average PSNR values of 15 test images using six different blur kernels. When the anchor point and the initial point coincide, GraRED-HP3 achieves the best performance on some IR tasks. Moreover, by dynamically updating the anchor points, restarted HPPP further boosts the performance of GraRED-HP3.
@@
Additionally, we trained a FNE-DnCNN ($D_{1}$) using real spectral normalization [ryu2019plug], which is employed in RED and RED-PRO. We also compared our results with state-of-the-art methods using advanced denoisers, such as DRUNet ($D_{2}$[zhang2021plug] and GS ($D_{3}$[hurault2022gradient, hurault2022proximal]. The PnP-DYS method [There-Operator-Splitting2024] uses the Davis-Yin splitting (DYS) algorithm. For GraRED-HP3, the anchor point is the same as the initial point. The default parameters are presented in Table 4. As shown in Table 6, Restarted HPPP with advanced denoisers achieves competitive or even superior performance compared to state-of-the-art methods. In Figure 8, we compare GraRED-P3 ($\lambda_{k}=0.1$) and GraRED-HP3 ($\mathbf{a}=\mathbf{u}^{0},\mu_{k}=\frac{1}{k+1}$). The results demonstrate that GraRED-HP3 achieves better recovery performance and exhibits a faster fixed-point convergence rate than GraRED-P3.
Table 3: Recovery results are obtained by different algorithms with trainable nonlinear reaction-diffusion (TNRD) denoiser [Chen2017]. The best two results are highlighted in red and blue, respectively.
@@
RED [Romano2017] 27.19 26.00 26.46 29.21 30.16 32.41 28.46 28.36 29.58
RED-PRO [cohen2021regularization] 27.43 26.05 26.39 28.99 29.98 32.41 28.25 28.21 29.38
DPIR [zhang2021plug] 28.47 27.02 27.63 30.24 30.89 33.56 29.62 30.43 30.54
PnP-DRS [hurault2022proximal] 29.79 26.92 27.46 29.70 29.59 33.25 29.94 30.66 29.46
PnP-PGD [hurault2022gradient] 27.18 26.28 26.73 29.17 29.14 32.84 29.14 29.87 30.37
PnP-DYS [There-Operator-Splitting2024] 27.22 26.27 26.62 28.97 29.04 32.87 29.06 29.88 28.68
@@ 4.5 Image inpainting
In this section, we use the proposed algorithm to solve TV-regularized image inpainting problems and compare their numerical results with CP [Chambolle2011], PPP [Bredies2022], and HPPP algorithms. The discrete image inpainting model is
@@
Both GraRED-P3 and GraRED-HP3 use DnCNN [ryu2019plug], other parameters are used:
@@ 5 Conclusions
This paper introduces a Halpern-type PPP (HPPP) algorithm, which has the advantage of strong convergence over PPP. In addition, GraRED-HP3 is proposed by integrating HPPP with denoiser priors for IR problems, which is an accelerated PnP-ADMM. Advantages of HPPP are demonstrated through several toy examples, and GraRED-HP3 achieves state-of-the-art performance on IR experiments. In the future, we plan to extend the definition of $\mathcal{M}$-monotonicity to $\mathcal{M}$-comonotonicity for nonconvex case [Bauschke2021].
@@ Acknowledgments
The authors thank Prof. Defeng Sun for bringing the interesting work [Sun2024] to our attention and Mr. Feng Xue for his insightful discussions. We sincerely appreciate Associate Editor Michael Elad and the anonymous referees for their thorough review and valuable comments, which have substantially improved this paper.
@@ Appendix A Boundedness
First, we study the boundedness and asymptotic regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by (5). To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce an important lemma from [2002Iterative, Lemma 2.5].
@@
Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ [Bredies2022, Proposition 2.3]. Denote
@@
First, we prove $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ is bounded. For any $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$, we have
@@ Appendix B $\mathcal{M}$-projection
As mentioned in [Bauschke2023, BUI2020124315], we can introduce the following notion of $\mathcal{M}$-projection.